OpenStudy (anonymous):

AB←→ and BC←→ form a right angle at point B. If A = (-3, -1) and B = (4, 4), what is the equation of BC←→?

OpenStudy (anonymous):

x + 3y = 16 2x + y = 12 -7x − 5y = -48 7x − 5y = 48

OpenStudy (anonymous):

@jayzdd

OpenStudy (anonymous):

@MrProphecyXII

OpenStudy (anonymous):

no need to answer it i need help on a different question

OpenStudy (anonymous):

no need to answer it i need help on a different question

OpenStudy (anonymous):

no need to answer it i need help on a different question

OpenStudy (anonymous):

no need to answer it i need help on a different question

OpenStudy (anonymous):

no need to answer it i need help on a different question

OpenStudy (anonymous):

no need to answer it i need help on a different question

OpenStudy (anonymous):

no need to answer it i need help on a different question

OpenStudy (anonymous):

no need to answer it i need help on a different question

OpenStudy (anonymous):

no need to answer it i need help on a different question

OpenStudy (anonymous):

no need to answer it i need help on a different question

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Ok,what is the other question?

OpenStudy (anonymous):

The transportation department has to put a railing along the outer curve of this road. The road is 20 feet wide. What is the length of the outer curve to the nearest foot? Use the value = 3.14. http://contentlaunch.ple.platoweb.com/testimagedb/5c/5ce075976004277a80bfbc0d167f00d2

OpenStudy (anonymous):

The transportation department has to put a railing along the outer curve of this road. The road is 20 feet wide. What is the length of the outer curve to the nearest foot? Use the value = 3.14. http://contentlaunch.ple.platoweb.com/testimagedb/5c/5ce075976004277a80bfbc0d167f00d2

OpenStudy (anonymous):

The transportation department has to put a railing along the outer curve of this road. The road is 20 feet wide. What is the length of the outer curve to the nearest foot? Use the value = 3.14. http://contentlaunch.ple.platoweb.com/testimagedb/5c/5ce075976004277a80bfbc0d167f00d2

OpenStudy (anonymous):

The transportation department has to put a railing along the outer curve of this road. The road is 20 feet wide. What is the length of the outer curve to the nearest foot? Use the value = 3.14. http://contentlaunch.ple.platoweb.com/testimagedb/5c/5ce075976004277a80bfbc0d167f00d2

OpenStudy (anonymous):

The transportation department has to put a railing along the outer curve of this road. The road is 20 feet wide. What is the length of the outer curve to the nearest foot? Use the value = 3.14. http://contentlaunch.ple.platoweb.com/testimagedb/5c/5ce075976004277a80bfbc0d167f00d2

OpenStudy (mathstudent55):

I can't open that file.

OpenStudy (anonymous):

sorry for the spam it wasnt loading